# Discrete time crystals in the presence of non-Markovian dynamics

### v1

@@ I Introduction
Many-body quantum systems driven out of equilibrium present exciting fields of research, owing to the different non-trivial behaviors exhibited by such systems, such as Kibble-Zurek mechanism [1, 2], dynamical localization [3, 4], and time-translational symmetry breaking [5, 6], to name a few. Recent developments in quantum simulators have enabled researchers to experimentally study the behavior of many-body systems driven out of equilibrium as well [7, 8]. Consequently, theoretical and experimental research on dynamics of many-body systems have received a lot of attention in the recent years. In particular, time-translational symmetry breaking in the form of continuous [9, 10, 11, 12, 13] or discrete [14, 15, 16] time crystals have received wide interest from the community in the last decade, and also realised experimentally [17, 18, 19, 20]
@@
Discrete time crystals (DTCs) are associated with spontaneous symmetry breaking in time, and are formed in periodically driven many-body quantum systems. In closed quantum systems DTCs have been studied widely, both theoretically [21, 14, 22, 23, 15] as well as experimentally [7], in many-body localized systems. More recently, DTCs have been shown to exist in periodically driven clean quantum systems as well [24, 25], and also in the presence of dissipation [26, 27, 28]. Such dissipative DTCs in periodically driven many-body open quantum systems can be viewed as quantum engines, wherein a part of the energy supplied through periodic drive flows to a cold bath, while the rest is obtained as output work [29, 6]. Analogous to different phases of matter, DTCs are robust to small perturbations. However, large perturbations, such as errors in the Hamiltonian [30], or strong rates of dissipation [27], can destroy a DTC. Consequently, finding scenarios which can result in stable DTCs is a fundamental question in the field of time-translational symmetry breaking in many-body systems. Furthermore, open questions remain regarding the existence and behavior of DTCs for different types of dissipative dynamics. Here we address the above two crucial issues by focusing on a periodically modulated many-body quantum system in the presence of non-Markovian dynamics. We show that suitably controlled non-Markovian dynamics can result in the generation of DTCs as well as other dynamical phases.
@@
Dissipative dynamics can be classified as Markovian or non-Markovian depending on the absence or presence of memory effects, respectively [31]. Markovian dynamics can be described by time-independent Lindblad superoperators, such that the system approaches a long-time steady-state monotonically. On the other hand, non-Markovian dynamics are associated with memory effects, which may result in the system moving away from the steady state for some intervals of time [32, 33, 34, 35].
@@
In this work, we focus on an open Dicke model, and present results in the Markovian limit, following Ref. [30]. In case of Markovian dynamics, DTC phase obtained for weak disspation is replaced by a time-independent steady-state (TISS) for intermediate or higher rates of dissipation [27]. In order to study the dynamics in the non-Markovian regime, we focus on a damped Jaynes-Cummings-like dissipator with tunable parameters, which allows one to traverse between the Markovian and the non-Markovian regimes [31]. We show that in contrast to the Markovian regime, the non-Markovian regime can be associated with a stable DTC for a wide range of parameter values, thereby significantly expanding the regime of parameters allowing the existence of time translational symmetry breaking. We present phase diagrams for the Markovian and non-Markovian regimes; striking differences emerge in the response of a system in the presence of Markovian and non-Markovian dissipation, for intermediate strengths of the dissipator coefficient. This presents the intriguing possibility of using dissipative DTC as a probe for detecting non-Markovianity at long times. Furthermore, we show that the DTC phase persists in the non-Markovian regime even in the presence of a random noise in the dissipator.
@@ II.1 Dissipative discrete time crystals
In contrast to Markovian dynamics, $\kappa(t)$ can assume negative values for some time intervals in the case of non-Markovian dynamics [32]. This results in the non-Markovian regime being associated with information backflow, such that a negative $\kappa(t)$ may drive $\mathcal{S}$ away from the long-time steady state for some time-intervals [35]. The distinct properties of Markovian and non-Markovian dynamics raises questions regarding the existence and characteristics of DTCs in the non-Markovian regime. Below we focus on the specific setup of a Dicke model in the presence of a Jaynes-Cummings-like dissipation, to show that indeed, the behavior of DTC in the non-Markovian regime can be significantly different from that seen in the Markovian regime.
@@ II.2 Modulated open Dicke model
The setup considered here posseses a $\mathcal{Z}_{2}$ symmetry; the Hamiltonian (2) commutes with the parity operator $P=e^{i\pi(\hat{a}^{\dagger}\hat{a}+\hat{J_{z}}+N/2)}$, such that $\hat{H}(\lambda)$ remains invariant under the transformation $\hat{a}\rightarrow-\hat{a},$ and $\hat{J_{x}}\rightarrow-\hat{J_{x}}$. However, in the thermodynamic limit $N\to\infty$, the $\mathcal{Z}_{2}$ symmetry is spontaneously broken leading to superradiant phase transition at the critical value of $\lambda=\lambda_{c}=\frac{1}{2}\sqrt{(\omega_{0}/\omega)(\omega^{2}+\kappa_{0}^{2}/4)}$ [37, 30].
In order to study the possibility of time-translational symmetry breaking in the presence of non-Markovian dynamics, we now consider a phenomenological model, wherein the above atom-photon setup (see Eq. (2)) evolves in the presence of a dissipative environment, following the master equation (1) [31]. In analogy with the damped Jaynes-Cummings model, here we focus on a time dependent photon decay rate $\kappa(t)$ given by
${\kappa(t)}=\begin{cases}\frac{2m\kappa_{0}\sinh(td/2)}{d\cosh(td/2)+m\sinh(td/2)}&\text{$|\kappa(t)|<\kappa_{max}$}\\ \kappa_{max}&\text{ $|\kappa(t)|\geq\kappa_{max}$},\\ \end{cases}$ (3)
where $d=\sqrt{m^{2}-2m\kappa_{0}}$ and $\mathcal{L}[\hat{a}]\hat{\rho}=\hat{a}\hat{\rho}\hat{a}^{\dagger}-\frac{1}{2}\{{\hat{a}^{\dagger}\hat{a},\hat{\rho}}\}$ (see Eq. (1)). Here $\kappa_{max}>\kappa_{0}$ is a parameter which can be tuned to control the maximum possible rate of dissipation. In case of the damped Jaynes-Cummings model, $\kappa_{max}\to\infty$, $m$ denotes the spectral width of the bath, while $\kappa_{0}$ is related to the system-bath coupling strength [31].
The above form of $\kappa(t)$ (Eq. (3)) allows us to tune between the Markovian regime ($\kappa_{0}<m/2$) in which case $\kappa(t)>0~{}\forall~{}t$, and the non-Markovian regime ($\kappa_{0}>m/2$), in which case $\kappa(t)$ assumes an oscillatory form realized by replacing $\sinh$ ($\cosh$) by $\sin$ ($\cos$) in Eq. (3), and can can take negative values for some time intervals [31, 35].
Following Ref. [30], one can use Eq. (1) to arrive at time-evolution equations for the scaled variables $x=\langle\hat{a}+\hat{a}^{\dagger}\rangle/\sqrt{2N\omega}$, $p=i\langle\hat{a}-\hat{a}^{\dagger}\rangle/\sqrt{2N\omega}$, and $\mathbb{j}=(j_{x},j_{y},j_{z})$ with $j_{\mu}=\langle\hat{J_{\mu}}\rangle/N$, in thermodynamic limit of $N\gg 1$ (see Appendix).
@@ II.2.1 Markovian regime
We start by focusing on the extreme Markovian limit of $m\gg\kappa_{0},~{}m\to\infty$, in which case $\kappa(t)$ reduces to the time-independent form $\kappa(t)\approx\kappa_{0}$ for all times. For a time-independent $\lambda_{t}=\lambda_{0}$, the above setup has two-symmetry broken steady states $\rho_{ss}$ and $\rho_{ss}^{\prime}$ (see Appendix) [38, 30]
@@
where, $T=\frac{2\pi}{\omega_{T}}$ , $\lambda_{0}>\lambda_{c}$ and $\omega_{T}=d/2$ [30]. In the resonant case $\omega=\omega_{0}$ and in absence of dissipation ($\kappa_{0}=0$), the above form of $\lambda_{t}$ results in the parity operation $\hat{P}=\exp[-i(T/2)H(0)+i\pi N/2]$ during the second half period $T/2\leq t<T$, up to a global phase which we can ignore.
@@
Figure 1: The mean magnetizations $j_{x}$, $j_{z}$, $j_{z}$ and the corresponding Bloch sphere representations are shown for the Markovian with $m\to\infty$ (a, b, c, d) and non-Markovian (e, f) regimes. We get a DTC phase for (a) small $\kappa_{0}=0.05$, represented by two dots in the (b) Bloch sphere. In contrast, a TISS is obtained for large $\kappa_{0}$ (c, d) in the Markovian regime. The DTC phase is preserved for intermediate values of $\kappa_{0}$ in the non-Markovian regime (e, f). Here $\omega_{T}=1$, $T=2\pi$, $\epsilon=0.02$, $\lambda_{0}=1$, a) $\kappa_{0}=0.05$, c) $\kappa_{0}=2.7$ c) $\omega_{T}=d/2$, $\kappa_{0}=2.7$ and $m=\kappa_{0}/4$, $T=2\pi/\omega_{T}$.
@@ II.2.2 Non-Markovian regime
In this section, we focus on the non-Markovian regime, obtained for $\kappa_{0}>m/2$. As discussed above, this regime corresponds to a periodically varying $\kappa(t)$ with $\kappa(t)<0$ for some time intervals (see Eq. (3)), which eventually results in the so called information backflow in the system [39].
In contrast to the behavior reported for Markovian dynamics [30] where DTC phase is present only for small $\kappa_{0}$ (see Figs. 1a - 1d), numerical analysis shows that non-Markovian dynamics makes DTC more robust against $\kappa_{0}$, as signified by the presence of time-crystalline order for intermediate values of $\kappa_{0}$ and $\kappa_{max}$; this is verified both for the Markovian limit with time-independent $\kappa(t)=\kappa_{0}$, obtained for $m/\kappa_{0}\to\infty$ (see Fig. 1), as well as for finite values of $m/\kappa_{0}$ (see Figs. 2 and 3). This robustness of the DTC phase w.r.t. $\kappa_{0}$ may be attributed to the periodically varying $\kappa(t)$ in the non-Markovian regime, even at long times [31, 40]. However, we note that the DTC phase is replaced by a TISS or a thermal phase for large $\kappa_{max}$ (see Appendix).
Figure 2: Stroboscopic dynamics (a,b,c,d), of the atomic average angular momentum $j_{\mu}$ for atom-photon coupling $\lambda=1$ and time dependent photon-loss rate $\kappa(t)$. As $\epsilon$ is varied, various dynamical phases emerges over long periods of time. Only the last 30 periods is shown in the figure. Last 200 periods (blue) is projected onto the bloch sphere. Here $\kappa_{0}=2.7$ and $\kappa_{max}=5$.
Figure 3: Phase diagram representing TISS, DTC, other dynamical phases (DP) and thermal phase (TP) as functions of $m$ and and $\kappa_{0}=2.7a)\kappa_{max}=5,b)\kappa_{max}=3$ , the dashed line at $m=2\kappa_{0}$, represents the transition line from NM to M regime.
@@
In order to have a deeper understanding of the effect of non-Markovian dynamics on the presence of DTC, we study the behavior of the system for different values of $\epsilon$ and $m/\kappa_{0}$. The results are summarized in the phase diagram Fig. 3. We navigate between the Markovian ($m/\kappa_{0}>2$) and the non-Markovian ($m/\kappa_{0}<2$) regimes by varying $m$, for a constant $\kappa_{0}$. Interestingly, as seen in Fig. 3, one can clearly distinguish the Markovian and non-Markovian regimes from the stark difference in behavior of the system across the transition. For a large enough $\kappa_{0}$ ($\kappa_{0}\gg T^{-1}$), the system approaches a TISS in the Markovian regime. However, varied dynamical phases, including DTC, appear on undergoing transition to the non-Markovian regime at $m=2\kappa_{0}$ (highlighted by a dashed black line in Fig. 3). Interestingly, the DTC phase shows the most robustness w.r.t. $\epsilon$ close to the Markovian to non-Markovian transition line $m=2\kappa_{0}$. This might be owing to small values of the parameter $d$ for $m\to 2\kappa_{0}^{-}$, which results in a slowly varying $\kappa(t)$, and an effectively longer time scales $\sim~{}d^{-1}$. As we move away from the transition line (i.e., $m\ll 2\kappa_{0}$), we find rich dynamical phases as we vary $\epsilon$. However, we note that this stark difference between Markovian and non-Markovian regimes vanishes for small $\kappa_{0}$, in which case varied dynamical phases may exist for different values of $\epsilon$ in the Markovian regime as well.

### v2

@@ I Introduction
Many-body quantum systems driven out of equilibrium present exciting fields of research [1], owing to the different non-trivial behaviors exhibited by such systems, such as Kibble-Zurek mechanism [2, 3], dynamical localization [4, 5], and time-translational symmetry breaking [6, 7], to name a few. Recent developments in quantum simulators have enabled researchers to experimentally study the behavior of many-body systems driven out of equilibrium as well [8, 9]. Consequently, theoretical and experimental research on dynamics of many-body systems have received a lot of attention in the recent years. In particular, time-translational symmetry breaking in the form of continuous [10, 11, 12, 13, 14, 15, 16] or discrete [7, 17, 18, 19] time crystals have received wide interest from the community in the last decade, and also realised experimentally [8, 20, 21, 22]
@@
Discrete time crystals (DTCs) are associated with spontaneous symmetry breaking in time, and are formed in periodically driven many-body quantum systems. In closed quantum systems DTCs have been studied widely, both theoretically [7, 23, 17, 24, 25] as well as experimentally [8], in many-body localized systems. More recently, DTCs have been shown to exist in periodically driven clean quantum systems as well [26, 27], and also in the presence of dissipation [28, 29, 30]. Such dissipative DTCs in periodically driven many-body open quantum systems can be viewed as quantum engines, wherein a part of the energy supplied through periodic drive flows to a cold bath, while the rest is obtained as output work [31, 7]. Analogous to different phases of matter, DTCs are robust to small perturbations. However, large perturbations, such as errors in the Hamiltonian [32], or strong rates of dissipation [29, 33], can destroy a DTC. Consequently, finding scenarios which can result in stable DTCs is a fundamental question in the field of time-translational symmetry breaking in many-body systems. Furthermore, open questions remain regarding the existence and behavior of DTCs for different types of dissipative dynamics. Here we address the above two crucial issues by focusing on a periodically modulated many-body quantum system in the presence of non-Markovian dynamics. We show that suitably controlled non-Markovian dynamics can result in the generation of DTCs as well as other dynamical phases.
@@
Dissipative dynamics can be classified as Markovian or non-Markovian depending on the absence or presence of memory effects, respectively [34]. Markovian dynamics can be described by time-independent Lindblad superoperators, such that the system approaches a long-time steady-state monotonically. On the other hand, non-Markovian dynamics are associated with memory effects, which may result in the system moving away from the steady state for some intervals of time [35, 36, 37, 38, 39].
@@
In this work, we focus on an open Dicke model, and present results in the Markovian limit, following Ref.[32]. In case of Markovian dynamics, DTC phase obtained for weak dissipation is replaced by a time-independent steady-state (TISS) for intermediate or higher rates of dissipation [29]. In order to study the dynamics in the non-Markovian regime, we focus on a Jaynes-Cummings-like dissipator with tunable parameters, which allows one to traverse between the Markovian and the non-Markovian regimes [34]. We show that in contrast to the Markovian regime, the non-Markovian regime can be associated with a stable DTC for a wide range of parameter values, thereby significantly expanding the regime of parameters allowing the existence of time translational symmetry breaking. We present phase diagrams for the Markovian and non-Markovian regimes; striking differences emerge in the response of a system in the presence of Markovian and non-Markovian dissipation, for intermediate strengths of the dissipator coefficient. This presents the intriguing possibility of using dissipative DTC as a probe for detecting non-Markovianity at long times. Furthermore, we show that the DTC phase persists in the non-Markovian regime even in the presence of a random noise in the dissipator.
@@ II.1 Dissipative discrete time crystals
In contrast to Markovian dynamics, $\kappa(t)$ can assume negative values for some time intervals in the case of non-Markovian dynamics [35]. This results in the non-Markovian regime being associated with information back-flow, such that a negative $\kappa(t)$ may drive $\mathcal{S}$ away from the long-time steady state for some time-intervals [39]. The distinct properties of Markovian and non-Markovian dynamics raises questions regarding the existence and characteristics of DTCs in the non-Markovian regime. For example, in a non-Markovian dynamics with a continuously varying $\kappa(t)$, the dissipative dynamics slows down significantly close to $\kappa(t)\to 0$, which may be beneficial for the stabilization of a DTC phase. Below we focus on the specific setup of a Dicke model in the presence of a Jaynes-Cummings-like dissipation, to show that indeed, the behavior of DTC in the non-Markovian regime can be significantly different from that seen in the Markovian regime.
@@ II.2 Modulated open Dicke model
The setup considered here possesses a $\mathcal{Z}_{2}$ symmetry; the Hamiltonian (2) commutes with the parity operator $P=e^{i\pi(\hat{a}^{\dagger}\hat{a}+\hat{J_{z}}+N/2)}$, such that $\hat{H}(\lambda)$ remains invariant under the transformation $\hat{a}\rightarrow-\hat{a},$ and $\hat{J_{x}}\rightarrow-\hat{J_{x}}$. However, in the thermodynamic limit $N\to\infty$, the $\mathcal{Z}_{2}$ symmetry is spontaneously broken leading to superradiant phase transition at the critical value of $\lambda=\lambda_{c}=\frac{1}{2}\sqrt{(\omega_{0}/\omega)(\omega^{2}+\kappa_{0}^{2}/4)}$ [41, 32].
In order to study the possibility of time-translational symmetry breaking in the presence of non-Markovian dynamics, we now consider a phenomenological model, wherein the above atom-photon setup (see Eq. (2)) evolves in the presence of a dissipative environment, following the master equation (1). In analogy with the Jaynes-Cummings model describing the dynamics of a two-level system coupled to a bath comprising Bosonic field modes characterized by a Lorentzian spectral function [34], we consider here a $\kappa(t)$ given by
${\kappa(t)}=\begin{cases}\frac{2m\kappa_{0}\sinh(td/2)}{d\cosh(td/2)+m\sinh(td/2)}&\text{$|\kappa(t)|<\kappa_{\rm{max}}$}\\ \kappa_{\rm{max}}&\text{ $|\kappa(t)|\geq\kappa_{\rm{max}}$}.\\ \end{cases}$ (3)
Here $d=\sqrt{m^{2}-2m\kappa_{0}}$, $\mathcal{L}[\hat{a}]\hat{\rho}=\hat{a}\hat{\rho}\hat{a}^{\dagger}-\frac{1}{2}\{{\hat{a}^{\dagger}\hat{a},\hat{\rho}}\}$ (see Eq. (1)), and $\kappa_{\rm{max}}>\kappa_{0}$ is a parameter which can be tuned to control the maximum possible rate of dissipation. In case of the Jaynes-Cummings model, $\kappa_{\rm max}\to\infty$, $m$ denotes the spectral width of the bath, while $\kappa_{0}$ is related to the system-bath coupling strength [34].
The above form of $\kappa(t)$ (Eq. (3)) allows us to tune between the Markovian regime ($\kappa_{0}<m/2$) in which case $\kappa(t)>0~{}\forall~{}t$, and the non-Markovian regime ($\kappa_{0}>m/2$), in which case $\kappa(t)$ assumes an oscillatory form realized by replacing $\sinh$ ($\cosh$) by $\sin$ ($\cos$) in Eq. (3), and can can take negative values for some time intervals (see Sec. II.2.2) [34, 39].
Following Ref. [32], one can use Eq. (1) to arrive at time-evolution equations for the scaled variables $x=\langle\hat{a}+\hat{a}^{\dagger}\rangle/\sqrt{2N\omega}$, $p=i\langle\hat{a}-\hat{a}^{\dagger}\rangle/\sqrt{2N\omega}$, and $\mathbb{j}=(j_{x},j_{y},j_{z})$ with $j_{\mu}=\langle\hat{J_{\mu}}\rangle/N$, in thermodynamic limit of $N\gg 1$ (see Appendix A).
@@ II.2.1 Markovian regime
We start by focusing on the extreme Markovian limit of $m\gg\kappa_{0},~{}m\to\infty$, in which case $\kappa(t)$ reduces to the time-independent form $\kappa(t)\approx\kappa_{0}$ for all times. For a time-independent $\lambda_{t}=\lambda_{0}$, the above setup has two-symmetry broken steady states $\rho_{ss}$ and $\rho_{ss}^{\prime}$ (see Appendix A) [42, 32]
@@
where, $T=\frac{2\pi}{\omega_{T}}$ and $\lambda_{0}>\lambda_{c}$ [32]. In the resonant case $\omega=\omega_{0}$ and in absence of dissipation ($\kappa_{0}=0$), the above form of $\lambda_{t}$ results in the parity operation $\hat{P}=\exp[-i(T/2)H(0)+i\pi N/2]$ during the second half period $T/2\leq t<T$, up to a global phase which we can ignore.
@@
In this Markovian regime, the parameter $\kappa(t)$ changes with time for small $t$, finally reaching the value $\kappa(t)=2m\kappa_{0}/(d+m)$ for long times $t\gg d^{-1}$. However, our numerical analysis shows that the behavior of the system at long times ($t\gg T,d^{-1}$) is independent of $d$.
@@
Figure 1: The mean magnetizations $j_{x}$, $j_{y}$, $j_{z}$ and the corresponding Bloch sphere representations are shown for the Markovian with $m\to\infty$ (a, b, c, d) and non-Markovian (e, f) regimes. We get a DTC phase for (a) small $\kappa_{0}=0.05$, represented by two dots in the (b) Bloch sphere. In contrast, a TISS is obtained for large $\kappa_{0}$ (c, d) in the Markovian regime. The DTC phase is preserved for intermediate values of $\kappa_{0}$ in the non-Markovian regime (e, f). Here $\omega_{T}=1$, $T=2\pi$, $\epsilon=0.02$, $\lambda_{0}=1$, a) $\kappa_{0}=0.05$, c) $\kappa_{0}=2.7$ c) $\omega_{T}=d/2$, $\kappa_{0}=2.7$ and $m=\kappa_{0}/4$, $T=2\pi/\omega_{T}$.
@@ II.2.2 Non-Markovian regime
In this section, we focus on the non-Markovian regime, obtained for $\kappa_{0}>m/2$. In this case $\kappa(t)$ assumes the form (see Eq. (3))
${\kappa(t)}=\begin{cases}\frac{2m\kappa_{0}\sin(t|d|/2)}{|d|\cos(t|d|/2)+m\sin(t|d|/2)}&\text{$|\kappa(t)|<\kappa_{\rm{max}}$}\\ \kappa_{\rm{max}}&\text{ $|\kappa(t)|\geq\kappa_{\rm{max}}$}.\\ \end{cases}$ (5)
This regime corresponds to a periodically varying $\kappa(t)$ with a time period $T_{\rm NM}=4\pi/|d|$; $\kappa(t)<0$ for some time intervals, which eventually results in the so called information back-flow in the system [34, 43]. Here for simplicity we consider a periodically modulated $\lambda_{t}$ with time period $T=T_{NM}$ (see Eq. (4)).
In contrast to the behavior reported for Markovian dynamics [32] where DTC phase is present only for small $\kappa_{0}$ (see Figs. 1a - 1d), numerical analysis shows that non-Markovian dynamics makes DTC more robust against $\kappa_{0}$, as signified by the presence of time-crystalline order for intermediate values of $\kappa_{0}$ and $\kappa_{\rm max}$; this is verified both for the Markovian limit with time-independent $\kappa(t)=\kappa_{0}$, obtained for $m/\kappa_{0}\to\infty$ (see Fig. 1), as well as for finite values of $m/\kappa_{0}$ (see Figs. 2 and 3). This robustness of the DTC phase w.r.t. $\kappa_{0}$ may be attributed to the periodically varying $\kappa(t)$ in the non-Markovian regime, even at long times [34, 44]. However, we note that the DTC phase is replaced by a TISS or a thermal phase for large $\kappa_{\rm{max}}$ (see Appendix B).
Figure 2: Stroboscopic dynamics (a,b,c,d), of the atomic average angular momentum $j_{\mu}$ for atom-photon coupling $\lambda=1$ and time dependent photon-loss rate $\kappa(t)$. As $\epsilon$ is varied, various dynamical phases emerges over long periods of time. Only the last 30 periods is shown in the figure. Last 200 periods (blue) are projected onto the Bloch sphere. Here $\kappa_{0}=2.7$ and $\kappa_{\rm{max}}=5$.
Figure 3: Phase diagram representing TISS, DTC, other dynamical phases (DP) and thermal phase (TP) as functions of $m$ and and $\kappa_{0}=2.7~{}a)\kappa_{\rm{max}}=5,b)\kappa_{\rm{max}}=3$ , the dashed line at $m=2\kappa_{0}$, represents the transition line from NM to M regime.
@@
In order to have a deeper understanding of the effect of non-Markovian dynamics on the presence of DTC, we study the behavior of the system for different values of $\epsilon$ and $m/\kappa_{0}$. The results are summarized in the phase diagram Fig. 3. We navigate between the Markovian ($m/\kappa_{0}>2$) and the non-Markovian ($m/\kappa_{0}<2$) regimes by varying $m$, for a constant $\kappa_{0}$. Interestingly, as seen in Fig. 3, one can clearly distinguish the Markovian and non-Markovian regimes from the stark difference in behavior of the system across the transition. For a large enough $\kappa_{0}$ ($\kappa_{0}\gg T^{-1}$), the system approaches a TISS in the Markovian regime. However, varied dynamical phases, including DTC, appear on undergoing transition to the non-Markovian regime at $m=2\kappa_{0}$ (highlighted by a dashed line in Fig. 3). Interestingly, the DTC phase shows the most robustness w.r.t. $\epsilon$ close to the Markovian to non-Markovian transition line $m=2\kappa_{0}$. This might be owing to small values of the parameter $d$ for $m\to 2\kappa_{0}^{-}$, which results in a slowly varying $\kappa(t)$, and an effectively longer time scales $\sim~{}d^{-1}$. As we move away from the transition line (i.e., $m\ll 2\kappa_{0}$), we find rich dynamical phases, including DTC which shows period doubling (see Fig. 2 b), sextet which shows periodic behavior with a time period of $6T$ (see Fig. 2 c; the corresponding Bloch sphere shows $6$ dots as the system oscillates between $6$ stable solutions as we vary $\epsilon$), and limit cycle where the system oscillates between two periodic orbits (see Appendix C). Other dynamical phases characterized by different time periods are shown in Fig. 2 d) and in the red shaded regimes in the Fig. 3. In addition, we get thermal phases, characterized by irregular trajectories and spins randomly distributed on the Bloch sphere (see Fig. 2 a). However, we note that this stark difference between Markovian and non-Markovian regimes vanishes for small $\kappa_{0}$, in which case varied dynamical phases may exist for different values of $\epsilon$ in the Markovian regime as well.